X-intercept -3, Y-intercept -2.

Grade:

2 Answers

Chetan Mandayam Nayakar
12 years ago

x-intercept=a, y-intercept=b, (x/a)+(y/b)=1

xyz xz
12 years ago

eqn of line is

 x/3 + y/2 =1

= 2x+3y-6=0